3-D SUBMILLIMETER SPECTROSCOPY OF ASTROPHYSICAL ’WEEDS’ – CONTINUED ANALYSIS

\underline{SARAH M. FORTMAN}, IVAN R. MEDVEDEV, CHRISTOPHER F. NEESE, and FRANK C. DE LUCIA, Department of Physics, 191 W. Woodruff Ave., The Ohio State University, Columbus, OH 43210-1106 USA.

In this presentation we report on the latest enhancements of the temperature resolved 3-dimensional spectroscopic technique and the recent results obtained. We have continued the analysis in the 210-270 GHz and 560-650 GHz spectral regions of the astrophysical weeds CH_{3}OH, CH_{3}CN, CH_{3}CHO, CH_{3}OCH_{3}, C_{2}H_{3}CN, C_{2}H_{5}CN, CH_{3}OCHO and SO_{2}. Our analysis includes a detailed comparison of the intensities of the quantum mechanical assignments and the experimental intensities to determine the completeness of each molecule as a function of temperature.